OpenStudy (anonymous):

solve the system of equations by graphing. Then classify the system. x+y=6 x-y=-4

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Move the X to the other side to isolate the Y. Y = -x +6 -Y = -x -4 The Y can't be negative, so divide all terms by -1. Y =X +4. So your equations are Y=-X+6 and Y= X+4. Do you have a graphing calculator? If so, then plug in those equations and a graph will come up.

OpenStudy (anonymous):

i do not have a graphing calculator

OpenStudy (anonymous):

Ok, I just went on an online graphing calculator and plugged in the equations. Your graph would look like this: